# Adapter-Enhanced Semantic Prompting for Continual Learning

### v2

@@ Abstract
Continual learning (CL) is essential for enabling models to adapt to dynamic data streams, with the primary challenge being the mitigation of catastrophic forgetting of previously acquired knowledge. Recent advancements have highlighted the potential of prompt-based CL methodologies, which have shown promise in preserving the knowledge embedded within pre-trained models, thereby addressing catastrophic forgetting. However, these methodologies are constrained by their focus on learning visual prompts for individual tasks, resulting in suboptimal generalization to unseen categories. To enhance the model’s ability to retain knowledge from prior tasks while simultaneously improving generalization to new tasks, we introduce a novel, lightweight framework for CL termed Adapter-Enhanced Semantic Prompting (AESP). This framework leverages class semantic information to augment visual features, thereby establishing relational links among different categories. This not only fortifies the retention of prior knowledge but also facilitates adaptation to new tasks. To seamlessly integrate visual and semantic information, we have developed a lightweight Adapter-enhanced Vision Transformer (ViT) architecture specifically designed for feature adaptation. Furthermore, we have implemented a robust Query-Key matching strategy to select the optimal task prompt pair, thereby enhancing the accuracy of final predictions. Comprehensive experiments conducted across three benchmark continual learning datasets demonstrate that our proposed framework outperforms several state-of-the-art approaches, showcasing its efficacy and robustness.
Index Terms:
@@
I Introduction
Continual learning (CL) [1],[2], [3], [4], [5], [6] is a paradigm designed to empower neural network models to seamlessly learn from evolving data streams and adapt to new environments. Unlike traditional machine learning models [7, 8], which are trained on static datasets and remain unchanged post-training, CL facilitates the accumulation of existing knowledge while assimilating new information from successive tasks. This dynamic approach obviates the necessity of retaining all historical task data and retraining models from scratch [3, 9], thereby addressing pivotal challenges such as data privacy and memory resource constraints. The inherent advantages of CL have spurred significant research interest in recent years [10], [11], [12].
Despite the advantage of CL, it is beset by the phenomenon of catastrophic forgetting [13], [14, 15], wherein the acquisition of new knowledge results in the overwrite of prior knowledge, leading to substantial performance degradation on previously learned tasks. To mitigate this issue, several strategies [16], [17, 18] have been proposed, including the retention of representative samples from past tasks for knowledge replay. Upon the arrival of a new task, these retained samples are rehearsed in conjunction with the new data, effectively reducing the model’s propensity to forget. Nevertheless, this method introduces a memory burden and potential privacy risks. Alternatively, approaches such as [3, 19] advocate for the addition of new branches to the network for each incoming task. While this strategy aids in knowledge preservation, it engenders an increase in model size, thereby compromising inference speed and overall efficiency.
@@ I Introduction
In recent years, prompt-based methodologies [20], [21], [22], [23], [24] have emerged as promising approaches for addressing the challenges of continual learning (CL). As illustrated in Figure 1 (a), these techniques integrate learnable visual prompts into image tokens during the training phase. The primary objective is to adapt pre-trained Vision Transformer (ViT) models to acquire new knowledge across successive CL tasks, while simultaneously mitigating forgetting by maintaining the ViT parameters in a frozen state. Distinct from traditional methods, prompt-based approaches necessitate the training of fewer parameters, yet they achieve superior performance by harnessing the extensive pre-trained knowledge embedded within ViT models.
@@
Nevertheless, a significant limitation of these methods is their reliance on visual information alone, which can constrain the model’s capacity to generalize effectively, especially when operating with limited and diverse training data. This constraint often results in ambiguity between newly acquired and previously learned categories, thereby exacerbating issues related to catastrophic forgetting. To address this limitation, leveraging class semantic information derived from large language models (LLMs) offers a more generalized and adaptable framework across diverse tasks [24], [23]. Notably, models such as CLIP [25] facilitate the alignment of samples within the same class with their corresponding semantic information, thereby constructing a more generalized feature space. This alignment significantly enhances the model’s robustness when encountering unseen samples from the same class, as it encourages the model to focus on shared semantic features rather than relying solely on specific visual cues.
@@
In this work, we introduce a novel Adapter-Enhanced Semantic Prompting (AESP) framework, specifically designed to advance the field of continual learning (CL). Unlike prior methods that focus solely on learning visual prompts, our approach leverages semantic embeddings derived from BERT’s text encoder [26] as semantic prompts, thereby enhancing the generalization of features encoded by the pre-trained Vision Transformer (ViT) model. However, a significant discrepancy persists between the representation spaces of semantic and visual information. The inherent attention mechanisms within a pre-trained ViT may not adequately reconcile these distinct modalities. To address this challenge, we introduce compact, trainable adapter modules within the ViT architecture, which are engineered to effectively fuse visual and semantic information. As illustrated in Figure 1 (b), these adapters are integrated into each layer of the ViT, enabling the model to better accommodate semantic information and fostering a more robust interaction between visual and semantic feature spaces. The AESP framework not only augments the integration of multimodal features but also endows the learned knowledge with enhanced generalization capabilities, thereby bolstering the model’s adaptability to subsequent tasks. Moreover, to ensure the precise selection of task-specific prompts for feature adaptation, we develop a novel Prompt-Key Matching (PKM) method. This method synthesizes a triad of matching strategies, ensuring the selection of the most relevant prompts from the repository, thereby optimizing task-matching accuracy and efficiency. Extensive experimental evaluations across various incremental tasks demonstrate the efficacy of our proposed AESP framework in significantly improving model performance.
The main contributions are summarized as follows:
@@
Figure 2: Main architecture of AESP. The semantic prompts generated by a large language model and learnable visual prompts are incorporated to enrich the input information (Section III-C). An integrated Query-Key matching mechanism is introduced to select the relevant prompts (Section III-D). The selected prompts alongside the image tokens are then fed into a novel backbone network, which includes adapters that can adapt to varying depths of the backbone network for feature refinement and learning (Section III-E). The final output features are optimized using classification loss and the proposed cosine contrastive loss to guide the model’s fitting process (Section III-F).
II Related work
II-A Continual Learning
@@
Continual Learning (CL) is a pivotal subdomain within the realm of machine learning, focused on the development of models capable of sequentially acquiring new knowledge while preserving previously acquired information [2, 3], [27, 5], [6]. The core challenge in this domain is the propensity for the acquisition of new knowledge to interfere with prior learning, resulting in the phenomenon known as catastrophic forgetting [13, 28, 29]. To mitigate this issue, CL techniques have evolved to incorporate a variety of strategies aimed at balancing the retention of existing knowledge with the assimilation of new information. Regularization methods represent one such strategy, wherein critical parameters are preserved by imposing constraints, thereby forestalling substantial modifications that could overwrite prior learning [30], [31, 32]. Knowledge distillation is another prominent approach, frequently employed to facilitate a seamless transition between legacy and updated models, thereby maintaining knowledge consistency [33, 34], [35]. Parameter isolation strategies [36, 37], [38] offer a different solution by safeguarding existing knowledge through the freezing of parameters associated with past tasks, while allocating new parameters for upcoming tasks. However, this approach can engender increased model complexity and necessitate greater maintenance efforts. Rehearsal methods [39, 40], [41] bolster prior learning by retraining on examples from previous tasks, although these methods are often constrained by the need for additional memory and may pose data privacy issues. The ongoing challenges posed by these methods underscore the need for innovative rehearsal-free techniques to effectively alleviate catastrophic forgetting in continual learning systems.
II-B Prompt-based Continual Learning
@@
In contrast to traditional methods that necessitate training a model from scratch, recent advancements in continual learning have shifted towards the fine-tuning of pre-trained networks [42, 43], [44, 45]. This paradigm is increasingly favored for its superior adaptability and learning efficiency. Prompt-based fine-tuning has emerged as a prominent strategy within this domain. Approaches such as L2P [20] and DualPrompt [21] represent innovative methodologies that leverage learnable prompts to dynamically steer the model’s attention during both training and inference phases. These prompts function as modular instructions that are integrated into the model, thereby encapsulating acquired knowledge and ensuring its ready accessibility, which is crucial for maintaining performance across diverse tasks. The incorporation of key-query mechanisms, as exemplified by methods like CODAPrompt [46], further enhances the precision of prompt-based learning by enabling a more nuanced, context-aware application of prompts. This advancement substantially broadens the model’s capacity to manage a wide array of learning scenarios without compromising existing knowledge. Nevertheless, the model’s predictive accuracy can be compromised if irrelevant prompts are selected during the inference process. To address this issue, CPrompt [47] introduces a consistent prompting strategy aimed at enhancing the robustness of prompts, thereby minimizing potential interference and ensuring more reliable performance.
II-C Adapter-based Continual Learning
@@
Adapter architectures [48, 49] were initially introduced for the fine-tuning of large language models (LLMs), enabling adaptation to new tasks through the addition of a minimal set of learnable parameters, while preserving the integrity of the backbone network’s parameters. This approach facilitates substantial parameter sharing, thereby enhancing efficiency. Building upon these foundational strengths, SSIAT [50] extends the application of adapters to the realm of continual learning, implementing a prototype shift strategy to counteract catastrophic forgetting induced by adapter parameter updates. This methodology has facilitated efficient fine-tuning of pre-trained networks, yielding performance that surpasses that of prompt-based approaches. C-ADA [9] adopts an expandable parameter strategy for adapters, coupled with regularization techniques to steer the trajectory of parameter updates, thereby mitigating interference between newly acquired and previously established knowledge. EASE [51] introduces a novel concept by designing a distinct, lightweight adapter module for each new task, and employs a semantic-guided prototype complement strategy to synthesize new features for existing classes, further contributing to the reduction of catastrophic forgetting. In contrast to these existing methodologies, our proposed framework innovates by incorporating a semantic prompt to enhance the generalization capacity of image features, and by designing adapters that seamlessly integrate semantic information with visual data. This integration aims to render the features more adaptable to the demands of continual learning tasks, representing a significant advancement in the field.
III Method
III-A Overview
@@
III-B Preliminary
Continual learning (CL) endeavors to equip models with the capability to assimilate knowledge from sequentially presented tasks with evolving data distributions, while maintaining competence in previously encountered tasks. Consider a dataset $\mathcal{D}$ comprising $N_{c}$ different classes $\mathcal{D}=\left\{\mathcal{D}_{1},\dots,\mathcal{D}_{\mathcal{T}}\right\}$, which is partitioned into $\mathcal{T}=N_{c}/N_{inc}$ incremental tasks, where $N_{inc}$ signifies the number of classes introduced in each successive task phase. The dataset for the $t$-th task is denoted as $\mathcal{D}_{t}=\left\{(x_{i}^{t},y_{i}^{t})\right\}_{i=1}^{N_{t}}$, where $N_{t}$ is the number of samples, $x_{i}^{t}$ is the $i$-th image in the $t$-th task, and the $y_{i}^{t}$ is its corresponding label. The label $y_{i}\in Y(t)$, where $Y(t)$ represents the class set in task $t$. After the $t$-th session, all the classes learned are denoted as $\mathcal{Y}=Y(1)\cup Y(2)\cup\dots\cup Y(t)$. To emulate real-world data streams, we employ a sequential data-feeding protocol. Our investigation is particularly focused on Class-Incremental Learning (CIL) [52, 41, 5], a scenario wherein models are trained with explicit task boundaries but are required to perform inference without knowledge of the task identity.
We utilize a Vision Transformer (ViT) [53], pre-trained on the ImageNet dataset, as the foundational architecture of our network. This architecture consists of an embedding layer $f_{e}$ followed by $N_{b}$ Multi-Head Self-Attention Blocks (MHSA) $\{f_{b}\}_{b=1}^{N_{b}}$. For an input image $x\in\mathbb{R}^{H\times W\times C}$ from the dataset $\mathcal{D}_{t}$, the embedding layer initially converts it into a token sequence, represented as a matrix $\xi_{e}=f_{e}(x)\in\mathbb{R}^{L_{img}\times d}$, where $L_{img}$ is the number of tokens and $d$ is the embedding dimension. Subsequently, this token sequence is augmented with a learnable class token $[CLS]\in\mathbb{R}^{d}$ and through the series of MHSA blocks to yield the final image representation $\xi_{cls}\in\mathbb{R}^{d}$. This representation $\xi_{cls}$ then used to train task-specific classifiers $\Phi(\cdot;\phi)$, where $\phi\in\left\{\phi_{1},\phi_{2},\dots,\phi_{\mathcal{T}}\right\}$ represents the parameters of the classifiers corresponding to each training task.
III-C Multimodal Prompt
Distinct from traditional methodologies that solely focus on learning visual prompts for feature adaptation, our approach integrates semantic prompts to augment the generalization capabilities of image features. These semantic prompts are derived from class semantic information. Specifically, we construct a semantic description $R_{lan}$ as follows:

### v3

@@ Abstract
Continual learning (CL) is essential for enabling models to adapt to dynamic data streams, with the primary challenge being the mitigation of catastrophic forgetting of previously acquired knowledge. Recent advancements have highlighted the potential of prompt-based CL methodologies, which have shown promise in preserving the knowledge embedded within pre-trained models, thereby addressing catastrophic forgetting. However, these methodologies are constrained by their focus on learning visual prompts for individual tasks, resulting in suboptimal generalization to unseen categories.
To enhance the model’s ability to retain knowledge from prior tasks while simultaneously improving generalization to new tasks, we introduce a novel, lightweight framework for CL termed Adapter-Enhanced Semantic Prompting (AESP). This framework leverages class semantic information to augment visual features, thereby establishing relational links among different categories. This not only fortifies the retention of prior knowledge but also facilitates adaptation to new tasks. To seamlessly integrate visual and semantic information, we have developed a lightweight Adapter-enhanced Vision Transformer (ViT) architecture specifically designed for feature adaptation. Furthermore, we have implemented a robust Query-Key matching strategy to select the optimal task prompt pair, thereby enhancing the accuracy of final predictions. Comprehensive experiments conducted across three benchmark continual learning datasets demonstrate that our proposed framework outperforms several state-of-the-art approaches, showcasing its efficacy and robustness.
Keywords:
@@
1 Introduction
Continual learning (CL)  (Aljundi et al., 2018; Belouadah et al., 2021; Lange et al., 2022; Tian et al., 2024; Zhou et al., 2025; Wang et al., 2024) is a paradigm designed to empower neural network models to seamlessly learn from evolving data streams and adapt to new environments. Unlike traditional machine learning models  (Jordan and Mitchell, 2015; Mahesh, 2020), which are trained on static datasets and remain unchanged post-training, CL facilitates the accumulation of existing knowledge while assimilating new information from successive tasks. This dynamic approach obviates the necessity of retaining all historical task data and retraining models from scratch  (Lange et al., 2022; Gao et al., 2024a), thereby addressing pivotal challenges such as data privacy and memory resource constraints. The inherent advantages of CL have spurred significant research interest in recent years  (Lopez-Paz and Ranzato, 2017; Parisi et al., 2019; Ren et al., 2019).
Despite the advantage of CL, it is beset by the phenomenon of catastrophic forgetting  (McCloskey and Cohen, 1989; French, 1999; Kong et al., 2023), wherein the acquisition of new knowledge results in the overwrite of prior knowledge, leading to substantial performance degradation on previously learned tasks. To mitigate this issue, several strategies  (Hou et al., 2019; Tao et al., 2020) have been proposed, including the retention of representative samples from past tasks for knowledge replay. Upon the arrival of a new task, these retained samples are rehearsed in conjunction with the new data, effectively reducing the model’s propensity to forget. Nevertheless, this method introduces a memory burden and potential privacy risks. Alternatively, approaches such as  (Lange et al., 2022; Bojian et al., 2023) advocate for the addition of new branches to the network for each incoming task. While this strategy aids in knowledge preservation, it engenders an increase in model size, thereby compromising inference speed and overall efficiency.
@@ 1 Introduction
In recent years, prompt-based methodologies  (Wang et al., 2022b, a; Tang et al., 2023; Khan et al., 2023; Liu et al., 2024) have emerged as promising approaches for addressing the challenges of continual learning (CL). As illustrated in Figure  1 (a), these techniques integrate learnable visual prompts into image tokens during the training phase. The primary objective is to adapt pre-trained Vision Transformer (ViT) models to acquire new knowledge across successive CL tasks, while simultaneously mitigating forgetting by maintaining the ViT parameters in a frozen state. Distinct from traditional methods, prompt-based approaches necessitate the training of fewer parameters, yet they achieve superior performance by harnessing the extensive pre-trained knowledge embedded within ViT models.
@@
Nevertheless, a significant limitation of these methods is their reliance on visual information alone, which can constrain the model’s capacity to generalize effectively, especially when operating with limited and diverse training data. This constraint often results in ambiguity between newly acquired and previously learned categories, thereby exacerbating issues related to catastrophic forgetting. To address this limitation, leveraging class semantic information derived from large language models (LLMs) offers a more generalized and adaptable framework across diverse tasks  (Khan et al., 2023; Park et al., 2024). Notably, models such as CLIP  (Radford et al., 2021) facilitate the alignment of samples within the same class with their corresponding semantic information, thereby constructing a more generalized feature space. This alignment significantly enhances the model’s robustness when encountering unseen samples from the same class, as it encourages the model to focus on shared semantic features rather than relying solely on specific visual cues.
@@
In this work, we introduce a novel Adapter-Enhanced Semantic Prompting (AESP) framework, specifically designed to advance the field of continual learning (CL). Unlike prior methods that focus solely on learning visual prompts, our approach leverages semantic embeddings derived from BERT’s text encoder  (Devlin et al., 2019) as semantic prompts, thereby enhancing the generalization of features encoded by the pre-trained Vision Transformer (ViT) model. However, a significant discrepancy persists between the representation spaces of semantic and visual information. The inherent attention mechanisms within a pre-trained ViT may not adequately reconcile these distinct modalities. To address this challenge, we introduce compact, trainable adapter modules within the ViT architecture, which are engineered to effectively fuse visual and semantic information. As illustrated in Figure  1 (b), these adapters are integrated into each layer of the ViT, enabling the model to better accommodate semantic information and fostering a more robust interaction between visual and semantic feature spaces. The AESP framework not only augments the integration of multimodal features but also endows the learned knowledge with enhanced generalization capabilities, thereby bolstering the model’s adaptability to subsequent tasks. Moreover, to ensure the precise selection of task-specific prompts for feature adaptation, we develop a novel Prompt-Key Matching (PKM) method. This method synthesizes a triad of matching strategies, ensuring the selection of the most relevant prompts from the repository, thereby optimizing task-matching accuracy and efficiency. Extensive experimental evaluations across various incremental tasks demonstrate the efficacy of our proposed AESP framework in significantly improving model performance.
The main contributions of our work are summarized as follows:
@@
2 Related Work
@@
Continual Learning.   Continual Learning (CL) is a pivotal subdomain within the realm of machine learning, focused on the development of models capable of sequentially acquiring new knowledge while preserving previously acquired information  (Belouadah et al., 2021; Lange et al., 2022; Cotogni et al., 2025). The core challenge in this domain is the propensity for the acquisition of new knowledge to interfere with prior learning, resulting in the phenomenon known as catastrophic forgetting  (McCloskey and Cohen, 1989; Shi et al., 2021). To mitigate this issue, CL techniques have evolved to incorporate a variety of strategies aimed at balancing the retention of existing knowledge with the assimilation of new information. Regularization methods represent one such strategy, wherein critical parameters are preserved by imposing constraints, thereby forestalling substantial modifications that could overwrite prior learning  (Nusrat and Jang, 2018; Ahn et al., 2019; Zhang et al., 2020). Knowledge distillation is another prominent approach, frequently employed to facilitate a seamless transition between legacy and updated models, thereby maintaining knowledge consistency  (Li et al., 2022, 2023; Xuan et al., 2025). Parameter isolation strategies  (Schwarz et al., 2018; Zhang et al., 2023; Zhao et al., 2023) offer a different solution by safeguarding existing knowledge through the freezing of parameters associated with past tasks, while allocating new parameters for upcoming tasks. However, this approach can engender increased model complexity and necessitate greater maintenance efforts. Rehearsal methods  (Bang et al., 2021; Bonicelli et al., 2022; Gao et al., 2023) bolster prior learning by retraining on examples from previous tasks, although these methods are often constrained by the need for additional memory and may pose data privacy issues. The ongoing challenges posed by these methods underscore the need for innovative rehearsal-free techniques to effectively alleviate catastrophic forgetting in continual learning systems.
@@
Prompt-based Continual Learning.   In contrast to traditional methods that necessitate training a model from scratch, recent advancements in continual learning have shifted towards the fine-tuning of pre-trained networks  (Chen et al., 2021; Pourpanah et al., 2023; Liu et al., 2023; Pei et al., 2024). This paradigm is increasingly favored for its superior adaptability and learning efficiency. Prompt-based fine-tuning has emerged as a prominent strategy within this domain. Approaches such as L2P  (Wang et al., 2022b) and DualPrompt  (Wang et al., 2022a) represent innovative methodologies that leverage learnable prompts to dynamically steer the model’s attention during both training and inference phases. These prompts function as modular instructions that are integrated into the model, thereby encapsulating acquired knowledge and ensuring its ready accessibility, which is crucial for maintaining performance across diverse tasks. The incorporation of key-query mechanisms, as exemplified by methods like CODAPrompt  (Smith et al., 2023), further enhances the precision of prompt-based learning by enabling a more nuanced, context-aware application of prompts. This advancement substantially broadens the model’s capacity to manage a wide array of learning scenarios without compromising existing knowledge. Nevertheless, the model’s predictive accuracy can be compromised if irrelevant prompts are selected during the inference process. To address this issue, CPrompt  (Gao et al., 2024b) introduces a consistent prompting strategy aimed at enhancing the robustness of prompts, thereby minimizing potential interference and ensuring more reliable performance.
@@
Adapter-based Continual Learning.   Adapter architectures  (Houlsby et al., 2019; Chen et al., 2022) were initially introduced for the fine-tuning of large language models (LLMs), enabling adaptation to new tasks through the addition of a minimal set of learnable parameters, while preserving the integrity of the backbone network’s parameters. This approach facilitates substantial parameter sharing, thereby enhancing efficiency. Building upon these foundational strengths, SSIAT  (Tan et al., 2024) extends the application of adapters to the realm of continual learning, implementing a prototype shift strategy to counteract catastrophic forgetting induced by adapter parameter updates. This methodology has facilitated efficient fine-tuning of pre-trained networks, yielding performance that surpasses that of prompt-based approaches. C-ADA  (Gao et al., 2024a) adopts an expandable parameter strategy for adapters, coupled with regularization techniques to steer the trajectory of parameter updates, thereby mitigating interference between newly acquired and previously established knowledge. EASE  (Zhou et al., 2024a) introduces a novel concept by designing a distinct, lightweight adapter module for each new task, and employs a semantic-guided prototype complement strategy to synthesize new features for existing classes, further contributing to the reduction of catastrophic forgetting. In contrast to these existing methodologies, our proposed framework innovates by incorporating a semantic prompt to enhance the generalization capacity of image features, and by designing adapters that seamlessly integrate semantic information with visual data. This integration aims to render the features more adaptable to the demands of continual learning tasks, representing a significant advancement in the field.
Figure 2: Main architecture of AESP. The semantic prompts generated by a large language model and learnable visual prompts are incorporated to enrich the input information (Section 3.3). An integrated Query-Key matching mechanism is introduced to select the relevant prompts (Section 3.4). The selected prompts alongside the image tokens are then fed into a novel backbone network, which includes adapters that can adapt to varying depths of the backbone network for feature refinement and learning (Section 3.5). The final output features are optimized using classification loss and the proposed cosine contrastive loss to guide the model’s fitting process (Section 3.6).
3 Method
3.1 Overview
@@
3.2 Preliminary
Continual learning (CL) endeavors to equip models with the capability to assimilate knowledge from sequentially presented tasks with evolving data distributions, while maintaining competence in previously encountered tasks. Consider a dataset $\mathcal{D}$ comprising $N_{c}$ different classes $\mathcal{D}=\left\{\mathcal{D}_{1},\dots,\mathcal{D}_{\mathcal{T}}\right\}$, which is partitioned into $\mathcal{T}=N_{c}/N_{inc}$ incremental tasks, where $N_{inc}$ signifies the number of classes introduced in each successive task phase. The dataset for the $t$-th task is denoted as $\mathcal{D}_{t}=\left\{(x_{i}^{t},y_{i}^{t})\right\}_{i=1}^{N_{t}}$, where $N_{t}$ is the number of samples, $x_{i}^{t}$ is the $i$-th image in the $t$-th task, and the $y_{i}^{t}$ is its corresponding label. The label $y_{i}\in Y(t)$, where $Y(t)$ represents the class set in task $t$. After the $t$-th session, all the classes learned are denoted as $\mathcal{Y}=Y(1)\cup Y(2)\cup\dots\cup Y(t)$. To emulate real-world data streams, we employ a sequential data-feeding protocol. Our investigation is particularly focused on Class-Incremental Learning (CIL)  (Mittal et al., 2021; Gao et al., 2023; Zhou et al., 2024b), a scenario wherein models are trained with explicit task boundaries but are required to perform inference without knowledge of the task identity.
We utilize a Vision Transformer (ViT)  (Dosovitskiy et al., 2021), pre-trained on the ImageNet dataset, as the foundational architecture of our network. This architecture consists of an embedding layer $f_{e}$ followed by $N_{b}$ Multi-Head Self-Attention Blocks (MHSA) $\{f_{b}\}_{b=1}^{N_{b}}$. For an input image $x\in\mathbb{R}^{H\times W\times C}$ from the dataset $\mathcal{D}_{t}$, the embedding layer initially converts it into a token sequence, represented as a matrix $\xi_{e}=f_{e}(x)\in\mathbb{R}^{L_{img}\times d}$, where $L_{img}$ is the number of tokens and $d$ is the embedding dimension. Subsequently, this token sequence is augmented with a learnable class token $[CLS]\in\mathbb{R}^{d}$ and through the series of MHSA blocks to yield the final image representation $\xi_{cls}\in\mathbb{R}^{d}$. This representation $\xi_{cls}$ then used to train task-specific classifiers $\Phi(\cdot;\phi)$, where $\phi\in\left\{\phi_{1},\phi_{2},\dots,\phi_{\mathcal{T}}\right\}$ represents the parameters of the classifiers corresponding to each training task.
3.3 Multimodal Prompt